| Operation Summary for December 7, 2011 - December 9, 2011 |
|
Wednesday December 7 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with NuMI, and MiniBooNE, experiments taking beam. During the day shift, operators turned on the Muon Rings, Transport, Pretarget, and Prevault at 8:15 AM in preparation for studies. At 8:15 AM, Operations prepared Transfer Hall and Enclosure B for controlled access. Operators restored MI crate #55 due to a 453 card that lost its parameters at 1:30 PM. During the evening shift, Operations disabled Muon Rings, Transport, Pretarget, and Prevault at 4:51 PM. At 7:58 PM a 453 card in MI crate #55 was replaced, this is the same card that caused problems on the day shift. At 9:12 PM, KRF5 tripped off and experts were called in to repair it. No beam anywhere for the rest of the shift, experts continued work on KRF5. T hursday December 8 The midnight shift began with no beam, KRF5 work continued. During the midnight shift, operators restored beam to NuMI and MiniBooNE, once KRF5 was repaired at 1:16 AM. Operators turned on the Muon Rings, Transport, Pretarget, and Prevault at 6:27 AM in preparation for studies. During the day shift, Operations set up the TLG at 11:16 AM for Recycler Ring studies. At 2:00 PM approval was given for T-1015 and T-1017 to take beam in the MTest beamline. Operations established beam to MTest at 4:00 PM. During the evening shift, the Muon Ring studies ended at 7:57 PM and operators disabled Muon Rings, Transport, Pretarget, and Prevault. At 8:58 PM a Booster bulk power supply (B:CEP400) tripped, operators reset it. Operations disabled MTest at 10:00 PM. MIRF14 tripped at 10:35 PM, operators have bypassed it, experts will look at it in the morning. Fri day December 9 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, operators monitored NuMI and MiniBooNE. Operations Coordinator (OpsCo) Plans for Today and for the Weekend Replace B:CEP400, return MIRF14 to service, normal running for the weekend. |
|
More Information Comments and Suggestions |
| last modified 12/09/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |